# Uncertainty Quantification Metrics for Deep Regression

### v2

@@ III-C2 Energy Based Regression (EBR)
$Z(x_{i};\theta)$ is approximated by Monte Carlo sampling. Predictions are generated by performing multi-start gradient ascent to maximize $f(x_{i},\tilde{y};\theta)$ with respect to $\tilde{y}$.
IV Experiments and Results
@@
IV-A Synthetic Regression Datasets
In order to gauge the behavior of the different uncertainty metrics, we construct four simple synthetic regression datasets, each with a different source of uncertainty. These can be seen in Fig. 2. We will henceforth refer to these datasets by their names: homoscedastic, heteroscedastic, multimodal, and epistemic. The names, homoscedastic and heteroscedastic are in reference to the type of Gaussian noise applied to the generating function. Multimodal refers to the fact that data is generated from two separate generating functions. Finally, epistemic, refers to the epistemic uncertainty arising from a gap in the training data. Fundamentally, each dataset is a collection of input-output pairs $(x,y)$. Outputs $y$ are generated as follows:
@@ IV-A Synthetic Regression Datasets
Here, $\mathcal{N}(0,\sigma)$ refers to a Gaussian distribution with mean 0, and standard deviation $\sigma$. When we write $\pm$ in the definition of the multimodal dataset, we refer to a 50/50 chance of being + or -, which represents the two different modes. In the epistemic dataset, there is a gap in the training data for $x\in[0.35,0.65]$, which is not present in the test data. This means test data is out of distribution, which causes epistemic uncertainty. Domains for the inputs are $x\in[-1,1]$ for the homoscedastic and heteroscedastic datasets, and $x\in[0,1]$ for the multimodal and epistemic datasets.
IV-B Implementation Details
For each synthetic dataset, we train one EBR model, and one DE model. Both types of models were trained with a batch size of 128 using the Adam [31] optimizer. Fig. 3 shows the resulting log-likelihood functions from the models, along with their predictions on our test sets.
Energy Based Regression Model We implement EBR model as a 9-layer perceptron with ReLU activations and a hidden size of 256, and we train for 20 epochs with a fixed learning rate of $10^{-4}$.
Deep Ensemble All models in our ensemble are identical 5-layer perceptrons, with ReLU activations, and a hidden size of 256. We train each model in the ensemble for 20 epochs with a fixed learning rate of $10^{-3}$. In our experiments we use the predicted variance as $U(x)$, more choices are compared in Xiong et al.  [23].
@@ IV-B Implementation Details
Figure 3: Visualization of the predicted density on the test set for trained models. Contour plots: Log-likelihood output from each model. Yellow: the high-density region; Blue: the low-density region. Blue points: Predicted mean. Orange points: Ground truth.
IV-C Stability on varying test set sizes
All metrics require a test set for evaluation. Since a test set is, by necessity, a finite set of data points, each metric is an approximation. Therefore it is important to investigate how quickly each metric converges to its expected value. Additionally, each metric may be a biased approximation. We perform two experiments to test these two different aspects of stability.
First, in order to test convergence, we emulate a process of iteratively collecting points. We begin by generating a single large test dataset of size $2^{16}$. From this dataset, we randomly sample data points without replacement, which allows us to investigate how the approximation of each metric behaves as the number of data points increases. We evaluate each metric at test set sizes $2^{3}$, $2^{4}$, …, $2^{16}$. The resulting metrics from each subset are reported in Fig. 4 (a).
@@ IV-C Stability on varying test set sizes
For both of these experiments, we use our heteroscedastic dataset because we believe that it is the best candidate to highlight any stability issues, due to the varying noise levels present in the dataset.
@@
IV-D Metrics under different types of uncertainty
For each dataset with different types of uncertainty, we compute all metrics for both our models on the test sets. Additionally, we wish to investigate how different metrics behave given the true data-generating distribution. Using the true data generating distribution will, in expectation, minimize NLL and calibration error, and should provide a good point of comparison for all metrics. We summarize metrics of both models for each dataset in Tab. I, II, III and IV respectively.
TABLE I: All metrics for all models on Homoscedastic dataset.
@@ IV-D Metrics under different types of uncertainty
Ensemble 0.5915 0.0023 -0.8819
EBR 0.5707 0.0032 -0.8568
True dist. 0.5917 0.0003 -0.8965
TABLE II: Metrics for both models on Heteroscedastic dataset.
@@
Ensemble 0.2334 0.0001 -0.1091
EBR 0.2422 0.0001 -0.0990
True dist. 0.2305 0.0001 -0.1472
TABLE III: Metrics for both models on multimodal dataset.
@@
Ensemble 0.0071 0.0229 0.8098
EBR 0.5821 0.0018 -0.6535
True dist. 0.5180 0.0001 -0.7935
TABLE IV: Metrics for both models on epistemic dataset.
@@
Ensemble 0.6016 0.0145 36.4332
EBR 1.3888 0.0298 31.2425
True dist. 0.5454 0.0001 -1.5871
V Discussion
@@
V-A Stability on varying test set sizes
@@
From Fig. 4 (a) we can see that all metrics vary as the test set sizes increase and finally converge to the expectation. We can conclude that all metrics converge beyond a dataset size of $2^{10}$ (=1024). Hence, for most modern regression applications, where test datasets are larger than $\sim$1000 samples, stability should not be a cause for concern. Though the amplitudes of metrics vary from each other, arguably we can also conclude that calibration error is the most stable and AUSE converges second fastest.
@@
From Fig. 4 (b) we can see that the averages change for different dataset sizes. We will henceforth refer to this behavior as estimation bias of mean. Arguably, estimation bias is undesirable, since we expect the average of a metric to converge to the same value regardless of the dataset size. As such, a metric with significant bias of the mean should be considered unstable and its use discouraged. Overall, calibration error is the most stable and AUSE the second most in terms of bias of the mean. We can also conclude that no metric seems to exhibit any meaningful bias beyond a dataset size of $2^{6}$. On the other hand, there is still visible bias with dataset size $2^{5}$, where the total number of samples collected in 100 test sets is much larger than the previously suggested size $2^{10}$. This indicates that it is advantageous to collect one large test set compared to aggregating many smaller test sets.
@@
Together, these two plots suggest that the most stable metric is calibration error, followed by AUSE, then NLL, and finally Spearman correlation. Surprisingly, in both Fig. 4 (a) and (b), Spearman correlation has a large value with small test set sizes and gradually converges to zero as the test set size grows. This can be explained by a closer look at Fig. 3 (b): All points close to $x=0$ will have similar predicted uncertainty values, but their errors may be vastly different, which counteracts the correlation between error and uncertainty. Naturally, this counteracting effect becomes less impactful when samples are sparse. By extension, it is also less impactful in higher-dimensional spaces, and Spearman correlation has been used successfully on for example the rMD17 dataset [32].
V-B Metrics under different types of uncertainty
Results for Spearman correlation will not be displayed in the following experiments as it converges to zero.
@@ V-B Metrics under different types of uncertainty
Homoscedastic and Heteroscedastic datasets: we first look at the predicted density distribution in Fig. 3 (a) and (b) for each model. We can see that both models learn distributions close to the generating functions (see Sec. IV-A) for both datasets.
@@
In Tab. I and II, all metrics suggest that both models learn the homoscedastic and heteroscedastic distributions approximately equally well. Interestingly, AUSE seems to imply that the EBR model has learned a distribution that is better than the ground truth of the homoscedastic dataset.
Refer to caption     Refer to caption
@@
Figure 5: Sparsification plots from Deep Ensemble, for the Homoscedastic and heteroscedastic datasets.
@@
Looking at Fig. 5 (a), the sparsification curve is nearly horizontal, because the predicted variances are nearly uniform over the entire test set. It is hence impossible to learn a meaningful correlation between errors and uncertainty. The predicted variance can only partially explain the error, the remaining part is caused by the difference between the annotation and the predicted mean. Intuitively, we expect this scenario of a perfectly uniform variance to be rare in real-world applications. This also indicates that generating uncertainty measures from the predicted distribution may not be the optimal approach for sorting samples based on uncertainty. For the heteroscedastic dataset, the true distribution achieves the best NLL, and the DE model is slightly better than EBR model. This is because th DE model learns a sharper distribution than EBR, as can be seen in the contour plots of Fig. 3 (a). NLL, being a proper scoring rule, addresses calibration and sharpness simultaneously [33].
@@
Multimodal dataset: we first look at the predicted density distribution in Fig. 3 (c) for each model. The energy-based model successfully learns both modes. Our ensemble was never designed to output multimodal predictions, and therefore it entirely fails to capture the multimodality.
@@
Tab. III shows larger discrepancy between metrics. All metrics suggest that the energy-based model has learned a good distribution that is close to the ground truth distribution. AUSE seems to imply that the DE model has learned a near-perfect uncertainty predictor, even significantly better than the ground truth. Calibration error and NLL both successfully reflect the ensemble’s failure to learn both modes. This highlights the main difference between AUSE (and Spearman correlation) and calibration error/NLL: AUSE measures a correlation between uncertainty and the absolute error, while calibration error and NLL measure how well the learned distribution captures the data. Naturally, this raises the question of which type of measurement is most useful, which we will discuss in Sec. V-D.
@@
Epistemic dataset: In Fig. 3 (d), both models exhibit somewhat larger uncertainty in the OOD region. However, neither model predicts a distribution that assigns high probability to the test data. Visually, the predicted distributions of both models look extremely similar. The only obvious perceptible difference is the slight bend in the center of the distribution of the EBR model.
@@
In Tab. IV, AUSE seems to imply that the ensemble has learned a distribution that is almost as good as the ground truth, while it says the EBR model failed spectacularly. Both calibration error and NLL agree that both models failed to learn a good distribution, but they disagree on which model is worse. The contours in Fig. 3 (d) also shows the DE model gives sharper distribution. It is worth noting that there is an entire field of study of Out-of-Distribution detection, which offers better methods for handling epistemic uncertainty. Hence, we are only interested in whether or not each metric can capture our models’ failure to learn the underlying distribution.
@@
Looking at the AUSE scores of 0.6016 and 1.3888 for the ensemble and EBR models respectively, we can infer that this small bend reduces the correlation between error and uncertainty enough to (more than) double the AUSE score. Intuitively, we would expect these two similar distributions to get similar scores, which is exactly the case with both calibration error and NLL. This raises questions regarding the stability of AUSE under small model variations.
V-C Interpretability
Finally, an important aspect of metrics is their interpretability. Calibration error reaches the lower bound 0 when the predicted distributions are close to the ground truth. The upper bound 1 is when all the test samples have nearly zero values of cumulative density. While calibration error is highly interpretable, the same cannot be said for NLL and AUSE. With NLL and AUSE, we can compare UQ across models, however, the lower and upper bounds are unclear.

### v4

@@ 3.3.2 Energy Based Regression (EBR)
$Z(x_{i};\theta)$ is approximated by Monte Carlo sampling. Predictions are generated by performing multi-start gradient ascent to maximize $f(x_{i},\tilde{y};\theta)$ with respect to $\tilde{y}$. For uncertainty quantification, the raw value of the energy function is used.
4 Experiments and Results
@@
4.1 Synthetic Regression Datasets
In order to gauge the behavior of the different uncertainty metrics, we construct four simple synthetic regression datasets, each with a different source of uncertainty. These can be seen in Fig. 2. We will henceforth refer to these datasets by their names: homoscedastic, heteroscedastic, multimodal, and epistemic. The names homoscedastic and heteroscedastic are in reference to the type of Gaussian noise applied to the generating function. Multimodal refers to the fact that data is generated from two separate generating functions. Finally, epistemic, refers to the epistemic uncertainty arising from a gap in the training data. Fundamentally, each dataset is a collection of input-output pairs $(x,y)$. Outputs $y$ are generated as follows:
  • 1.
@@ 4.1 Synthetic Regression Datasets
Here, $\mathcal{N}(0,\sigma)$ refers to a Gaussian distribution with mean 0, and standard deviation $\sigma$. $\pm$ in the definition of the multimodal dataset refers to a equal chance of being + or -, which represents the two different modes. In the epistemic dataset, there is a gap in the training data for $x\in[0.35,0.65]$, which is not present in the test data. This means test data is OOD, which causes epistemic uncertainty. Domains for the inputs are $x\in[-1,1]$ for the homoscedastic and heteroscedastic datasets, and $x\in[0,1]$ for the multimodal and epistemic datasets.
4.2 Implementation Details
For each synthetic dataset, we train one EBR model, and one DE model. Both types of models were trained with a batch size of 128 using the Adam (Kingma & Ba, 2015) optimizer. The sizes of the training, validation, and testing datasets were respectively, 100000, 1000, and 1000 samples. Fig. 3 shows the resulting log-likelihood functions from the models, along with their predictions on our test sets.
Energy Based Regression We implement EBR as a 9-layer perceptron with ReLU activations and a hidden size of 256, which we train for 20 epochs with a fixed learning rate $10^{-4}$. While 9 layers may sound too deep for our simple toy datasets, our initial experiments showed clear signs of underfitting with shallower models.
Deep Ensemble All models in our ensemble are identical 5-layer perceptrons, with ReLU activations, and a hidden size of 256. In this case 5 layers is more than enough for the model to learn our toy datasets, and we choose this size simply to avoid a comparison between models of drastically different depth. Overfitting is not a concern since we can control the amount of data we generate. We train five models for 20 epochs each with a fixed learning rate of $10^{-3}$. In our experiments we use the predicted variance as $U(x)$, more choices are compared by Xiong et al. (2024).
@@ 4.2 Implementation Details
Figure 3: Visualization of the predicted density on the test set for trained models. Contour plots: Log-likelihood output from each model. Yellow: the high-density region; Blue: the low-density region. Blue points: Predicted mean. Orange points: Test set.
4.3 Stability on varying test set sizes
All metrics are estimates that are computed on a test set of finite size. It is thus important to investigate how quickly each metric converges to its expected value, and whether the estimates are biased for small test set sizes.
[…]
@@ 4.3 Stability on varying test set sizes
For both experiments, we use our heteroscedastic dataset because we believe it is the best candidate to highlight any stability issues, due to its varying noise levels. We apply the DE model in both experiments since its performance is similar to EBR model on the heteroscedastic dataset as shown in Fig. 3 (b) and the choice of model is agnostic.
@@
4.4 Metrics under different types of uncertainty
We compute all metrics for both our models on the test sets of each of the four datasets. As a reference, we compare the results to those computed on the data-generating distribution. Using the data generating distribution will, in expectation, minimize NLL and CE, and should provide a good point of comparison for all metrics. We summarize metrics of both models for each dataset in Tab. 1, 2, 3 and 4 respectively.
Table 1: Metrics for all models on Homoscedastic dataset.
@@ 4.4 Metrics under different types of uncertainty
DE 0.5915 0.0023 -0.8819 0.0444
EBR 0.5707 0.0032 -0.8568 -0.0264
True dist. 0.5917 0.0003 -0.8965 0.0076
Table 2: Metrics for both models on Heteroscedastic dataset.
@@
DE 0.2334 0.0001 -0.1091 0.0005
EBR 0.2422 0.0001 -0.0990 -0.0419
True dist. 0.2305 0.0001 -0.1472 0.0280
Table 3: Metrics for both models on multimodal dataset.
@@
DE 0.0071 0.0229 0.8098 -0.0420
EBR 0.5821 0.0018 -0.6535 0.0093
True dist. 0.5180 0.0001 -0.7935 0.0073
Table 4: Metrics for both models on epistemic dataset.
@@
DE 0.6016 0.0145 36.4332 0.0298
EBR 1.3888 0.0298 31.2425 0.0067
True dist. 0.5454 0.0001 -1.5871 -0.0194
4.5 Metrics for Real-world Applications
To further demonstrate the behaviour of the metrics, we apply them to the real-world problem of stereo disparity prediction. This is a regression task where the disparity of X-coordinates are to be predicted for corresponding pixels in the left and right frames of a stereo camera. We adopt the continuous disparity network(CDN) from Garg et al. (2020), and we use the Hinge-Wasserstein training loss (Xiong et al., 2024) (hinge value 0.005) as it improves both the regression and UQ performance. The model is trained on the large dataset Sceneflow (Mayer et al. (2016)), where dense ground truth disparity maps are provided. As this is real image data, there is already an unknown amount of aleatoric uncertainty in the test set. To this we add controllable amounts of uniform Gaussian noise, to both left and right frames during testing. Specifically, Gaussian noise with a set of different variances is added to the RGB image normalized by ImageNet mean and variance. (This means that the image variance can be expected to be around 1 in the noise free case.) We use End-Point-Error (EPE) and 1-Pixel Threshold Error (1PE), as in Garg et al. (2020) to quantify for regression performance. All relevant performance and uncertainty metrics are provided in Tab. 5.
Table 5: All metrics for stereo disparity prediction with CDN supervised by hinge-Wasserstein (hinge value 0.005). Variances are reported with 5 runs.
Noise scale EPE 1px NLL AUSE CE Spearman
0 1.01 ±$\pm 0$ 9.44 ±$\pm 0$ 0.000521±$\pm 3.6\times 10^{-6}$ 0.159 ±$\pm 0.0006$ 0.2663 ±$\pm 0$ 0.0012±$\pm 0.0004$
0.1 1.13 ±$\pm 0.00$ 10.97 ±$\pm 0.00$ 0.000590±$\pm 4.2\times 10^{-6}$ 0.161 ±$\pm 0.0014$ 0.2629 ±$\pm 0.0001$ 0.0010±$\pm 0.0006$
0.2 1.85 ±$\pm 0.00$ 15.91 ±$\pm 0.00$ 0.000735±$\pm 2.8\times 10^{-6}$ 0.128 ±$\pm 0.0013$ 0.2457 ±$\pm 0$ $\pm 0.0006$
0.3 2.81 ±$\pm 0.01$ 22.54 ±$\pm 0.01$ 0.000904±$\pm 1.2\times 10^{-6}$ 0.113 ±$\pm 0.0008$ 0.2210 ±$\pm 0$ 0.0024±$\pm 0.0005$
0.4 4.01 ±$\pm 0.01$ 29.81 ±$\pm 0.01$ 0.001061±$\pm 1.9\times 10^{-6}$ 0.103 ±$\pm 0.0008$ 0.1947 ±$\pm 0.0002$ 0.0026±$\pm 0.0006$
0.5 5.50 ±$\pm 0.01$ 37.32 ±$\pm 0.01$ 0.001212±$\pm 2.0\times 10^{-6}$ 0.099 ±$\pm 0.0008$ 0.1686 ±$\pm 0.0001$ 0.0040±$\pm 0.0757$
5 Discussion
@@
5.1 Stability on varying test set sizes
@@
From Fig. 4 (a) we can see that all metrics vary as the test set sizes increase and finally converge to the expectation. We can conclude that all metrics converge beyond a dataset size of $2^{10}$. Hence, for most modern regression applications, where test datasets are larger than $\sim$1000 samples, stability should not be a cause for concern. While our analysis is based only on a single dataset, we expect the convergence speed to be reasonably consistent across different datasets. Though the amplitudes of metrics vary from each other, arguably we can also conclude that CE is the most stable and AUSE converges second fastest.
@@
From Fig. 4 (b) we can see that the averages change for different dataset sizes. We will henceforth refer to this behavior as estimation bias of mean. Arguably, estimation bias is undesirable, since we expect the average of a metric to converge to the same value regardless of the dataset size. As such, a metric with significant bias of the mean should be considered unstable and its use discouraged. Overall, CE is the most stable and AUSE the second most in terms of bias of the mean. We can also conclude that no metric seems to exhibit any meaningful bias beyond a dataset size of $2^{6}$.
@@
Together, these two plots suggest that the most stable metric is CE, followed by AUSE, then NLL, and finally Spearman correlation. Surprisingly, in both Fig. 4 (a) and (b), Spearman correlation has large (absolute) values with small test set sizes and gradually converges to zero as the test set size grows. This can be explained by a closer look at Fig. 3 (b): All points close to $x=0$ will have similar predicted uncertainty values, but their errors may be vastly different, which counteracts the correlation between error and uncertainty. Naturally, this counteracting effect becomes less impactful when samples are sparse, and Spearman correlation has been used successfully on for example the rMD17 dataset (Christensen, 2020).
5.2 Metrics under different types of uncertainty
@@
Homoscedastic and Heteroscedastic datasets: In Fig. 3 (a) and (b) we can see that both models learn predictive distributions close to the generating functions (defined in Sec. 4.1).
@@
In Tab. 1 and 2, all metrics except Spearman suggest that both models learn the homoscedastic and heteroscedastic distributions approximately equally well.
@@
Figure 5: Sparsification plot from Deep Ensemble and True Distribution for the homoscedastic and heteroscedastic datasets. $\alpha$ is the fraction of removed samples.
@@
In Fig. 5 (a), the sparsification curve is nearly horizontal, indicating that it is impossible to learn a meaningful correlation between errors and uncertainty. The predicted variance can only partially explain the error, the remaining part is caused by the difference between the annotation and the predicted mean. Intuitively, we expect this scenario of a perfectly uniform variance to be rare in real-world applications. This also indicates that generating uncertainty measures from the predicted parametric distribution may not be the optimal approach for sorting samples. Fig. 3 (a) and Fig. 5 (b) together further show this: even when the predictive and generating distributions are very well aligned, the AUSE value is not 0. For the heteroscedastic dataset, the true distribution achieves the best NLL, and the DE model is slightly better than EBR model. This is because the DE model learns a sharper distribution than EBR, as can be seen in the contour plots of Fig. 3 (a). NLL, as a proper scoring rule, addresses calibration and sharpness simultaneously (Gneiting & Raftery, 2007).
@@
Multimodal dataset: In Fig. 3 (c) we can see that energy-based model successfully learns to predict both modes, but as the ensemble is not designed to output multimodal predictions, it predicts the average of the modes, and a wide distribution.
@@
Tab. 3 shows larger discrepancy between metrics. All metrics suggest that the energy-based model has learned a good distribution that is close to the generating distribution. AUSE for DE is near prefect, which indicates good detection of when DE fails to predict multiple modes. CE and NLL instead indicate the ensemble’s failure to learn both modes with high values. This highlights the main difference between AUSE (and Spearman correlation) and CE/NLL: AUSE measures a correlation between uncertainty and the absolute error, while CE and NLL measure how well the learned distribution captures the data. This raises the question of which type of measurement is most useful, which we will discuss in Sec. 5.5.
@@
Epistemic dataset: In Fig. 3 (d), both models exhibit somewhat larger uncertainty in the OOD region. However, neither model predicts a distribution that assigns high probability to the test data. Visually, the predicted distributions of both models look similar. The only obvious perceptible difference is the slight bend in the center of the distribution of the EBR model.
@@
In Tab. 4, AUSE indicates that the ensemble predicts distribution that can find OOD samples almost as well as the generating distribution, while the distribution from the EBR model is much worse. Both CE and NLL agree that both models failed to learn a good distribution, but they disagree on which model is worse. The contours in Fig. 3 (d) also shows the DE model gives sharper distribution. It is worth noting that there is an entire field of study of OOD detection, which offers better methods for handling epistemic uncertainty. Here, we are instead interested in whether or not each metric can detect a failure to learn the underlying distribution.
@@
Looking at the AUSE scores of 0.6016 and 1.3888 for the ensemble and EBR models respectively, we can infer that this small bend reduces the correlation between error and uncertainty enough to (more than) double the AUSE score. Intuitively, we would expect these two similar distributions to get similar scores, which is exactly the case with both CE and NLL. This raises questions regarding the stability of AUSE under small model variations.
5.3 Real-world Applications
Tab. 5 summarize the results of the real-world stereo disparity experiment. As can be expected, regression performances drop as the noise scale increases. It is immediately clear that the variance of Spearman correlation is too large to be a reliable metric. Both AUSE and CE keeps decreasing as the noise scale increases, which indicates that the used Hinge–Wasserstein loss is able to quantify uncertainty under adversarial noise. Meanwhile, NLL keeps increasing, which means the sharpness of predictions get worse with higher noise levels. This is consistent with Xiong et al. (2024) which illustrates that a lack of sharpness is a side effect of the Hinge–Wasserstein loss.
5.4 Interpretability
CE has well defined bounds which makes it easy to interpret. With NLL and AUSE, this is more difficult. We can compare them across models, however, the lower and upper bounds depend on both model and data.
CE: CE has a lower bound 0 when the predicted and generating distributions are equal. The upper bound, however, is dependent on the number of intervals and the weights $w_{j}$.